# Understanding and Mitigating the Threat of Vec2Text to Dense Retrieval Systems

### v1

@@ 4 Understanding what impacts Vec2Text effectiveness
To maintain control over experimental parameters and prevent results from being influenced by factors such as different datasets and training parameters, we independently train the embedding models ourselves. This is done with the same experimental settings but utilizing different training or inference strategies. Specifically, we start by reproducing the DPR embedding model Karpukhin et al. (2020), following the training parameters outlined in the DPR paper and employing the Tevatron dense retriever training toolkit Gao et al. (2023) to train the DPR model on the NQ dataset, the same dataset used to train Vec2Text. After reproducing the DPR retrieval results, we then apply the same dense retriever training hyperparameters (e.g., batch size, learning rate, and number of hard negatives) to train other DPR variations with distinct training strategies, as detailed in the next subsections.
4.1 Distance Metric and Pooling Method
@@
Regarding reconstructibility, a comparison between DRP_cls_dot and DRP_cls_cos suggests that cosine similarity appears to enhance Vec2Text’s exact matching score. However, when comparing DRP_cls_cos and DRP_mean_cos, the reconstructibility of Vec2Text significantly increases. This surprising result suggests that mean pooling is a key factor for Vec2Text reconstructibility but does not necessarily contribute to improved retrieval effectiveness. We suspect that this might be due to mean pooling explicitly gathering all the token information into the embedding, whereas the CLS token embedding only implicitly learns the text information during DR training, making it less helpful for Vec2Text training. Consequently, practitioners in the field of DR should carefully consider avoiding the use of mean pooling to enhance the security of the embedding model against Vec2Text attacks.
@@
Table 3: Zero-shot retrieval and bottlenecked pre-training results. Subscripts denote statistically significantly better retrieval effectiveness with a T-test ($p<0.01$). Results show that even zero-shot retrievers are at risk to Vec2Text reconstruction. Bottleneck pre-training proved useful at improve retrieval effectiveness but greatly increased risk of reconstructibility.
Embedding Models Retrieval effectiveness Reconstructibility
Label Setting top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. BERT_cls_dot Zero-shot 0.101 0.153 0.288 0.467 37.034 0.671 0.042 0.968
[…]
4.2 Zero-shot Regime and Bottleneck Pre-training
@@
BERT_cls_dot and BERT_mean_cos use the same backbone pre-trained embedding model (BERT Kenton and Toutanova (2019)) and perform the exact same inference as DRP_cls_dot and DRP_mean_cos but without DPR training; they represent zero-shot retrievers. Their zero-shot retrieval effectiveness is significantly worse than fine-tuned DPR models (as might be expected for a zero-shot dense retriever). Mean pooling with cosine similarity exhibits higher zero-shot retrieval effectiveness than dot product with CLS token embeddings but still well below the trained DPR models from the previous section.
@@
We then tested the SimLM Wang et al. (2023) embedding model, which leverages a bottlenecked pre-training approach Gao and Callan (2021); Lu et al. (2021); Liu and Shao (2022); Shen et al. (2022); Chuang et al. (2022); Wu et al. (2022); Zhuang et al. (2023). Bottlenecked pre-training is a task designed to enhance the CLS token embedding. This is achieved by taking the embedding model’s outputted CLS token embedding and inputting it into a weaker decoder model. The pre-training objective is to enable the decoder model to perform the MASK language modelling task or, in the case of SimLM, the ELECTRA pre-training task Clark et al. (2019). Consequently, the model must learn to inject useful information into the CLS token embedding to assist the decoder in completing the task. In fact, this pre-training approach closely resembles Vec2Text training itsself, as both involve a generator that takes an embedding as input and attempts to reconstruct a piece of text. Previous studies on bottlenecked pre-training have shown improvements in DR retrieval effectiveness. However, we suspect that this pre-training approach may make the embedding particularly vulnerable to Vec2Text, raising potential privacy concerns.
@@
Table 4: Impact of embedding dimension and quantization. Both dimension reduction and product quantization prove to be simple and effective methods to mitigate Vec2Text reconstruction, without comprising retrieval effectiveness. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot (768) 768 None 61GB 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_dot (256) 256 None 21GB 0.731 0.784 0.855 0.910 25.060 0.565 0.035 0.881
c. DPR_cls_dot (PQ_768) 768 Product 16GB 0.749 0.801 0.862 0.914 1.916 0.146 0.000 0.748
d. DPR_cls_dot (PQ_256) 256 Product 5GB 0.740 0.796 0.864 0.912 2.305 0.159 0.000 0.767
4.3 Embedding Dimensionality & Quantization
@@
Regarding embedding dimensionality, we trained another DPR model with the exact settings as DPR_cls_dot, but adding a dense pooling layer on top to reduce the embedding dimension from 768 to 256. This variant is labeled as DPR_cls_dot (256). For embedding quantization, we applied product quantization (PQ) Jegou et al. (2010) to DPR_cls_dot. We experimented with two settings: one with the number of sub-vectors set to 768 and another with 256. The number of bits per sub-vector for both settings is set to 8 in both cases, resulting in embedding dimensions of 768 and 256. We label them DPR_cls_dot (PQ_768) and DPR_cls_dot (PQ_256) respectively.
The results presented in Table 4 indicate that dimensionality reduction and product quantization can significantly reduce the dense vector index sizes, as expected. Product quantization demonstrates robust retrieval effectiveness; when the sub-vector is set to 768, there is no decrease in top-k accuracy, and only a slight decrease in retrieval effectiveness when further reducing the dimensionality to 256. These findings align with previous studies Ma et al. (2021).
@@
Table 5: Impact of adding noise and customized embedding transformation. Noise injection was an effective reconstruction mitigation strategy. The proposed vector transformation mitigation strategy is easily applied and guaranteed to maintain retrieval effective while completely degrading reconstructibility. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot None 0.748 0.800 0.863 0.914 70.735 0.875 0.319 0.985
b. DPR_cls_dot_noise Noise injection 0.730 0.790 0.856 0.911 10.022 0.415 0.000 0.823
c. DPR_cls_dot_transform Vector transformation 0.748 0.800 0.863 0.914 1.178 0.065 0.000 0.443
5 Mitigation Strategies
5.1 Noise Injection
$\phi_{noisy}(x)=\phi(x)+\lambda\cdot\epsilon,\epsilon\sim\mathcal{N}(0,1),$ (1)
Table 6: Results of adding noise during Text2Text training. The target embedding model is $DPR\_cls\_dot$. Noise injection proved to be an effective mitigation strategy even when Vec2Text is actually trained with noisy embeddings.
@@ 5.1 Noise Injection
Vec2Text Original embeddings 70.735 0.875 0.319 0.985 10.022 0.415 0.000 0.823
Vec2Text_noise Noisy embeddings 29.264 0.635 0.012 0.966 17.417 0.504 0.005 0.882
Figure 2: Impact on retrieval effectiveness and reconstructibility of different amounts of noise injection with different retrieval models. Larger $\lambda$ signifies more noise injection.
@@ 5.2 Embedding Transformation
6 A New Threat: Corpus Poisoning
In this section, we investigate whether Vec2Text poses a threat to existing DR systems through a corpus poisoning attack.
6.1 Existing Corpus Poisoning Attack Method
A corpus poisoning is an adversarial attack where a malicious user injects adversarial passages into a retrieval corpus with the intention of misleading DR models. This is a relatively new type of attack, and, to date, there is only one existing study conducted on this problem Zhong et al. (2023). Specifically, Zhong et al. (2023) proposed a gradient-based approach, inspired by the HotFlip method Ebrahimi et al. (2018); Wallace et al. (2019), which generates a small set of adversarial passages by perturbing discrete tokens in randomly initialized passages to maximize their similarity with a provided set of training queries. These adversarial passages are then inserted into the retrieval corpus, and the success of the attack is determined by the retrieval of these adversarial passages at a high rank in response to future unseen queries. These adversarial passages can be used to harm retrieval effectiveness and/or inject spam or content with misinformation into the search engine result list.
The current HotFlip-based corpus poisoning method has two drawbacks (that is, from an attacker perspective). Firstly, being a gradient-based method, it requires access to the embedding model weights, i.e., it is a white box attack. Consequently, attackers cannot employ this method to target DR systems built with closed-source embedding models, such as OpenAI embedding models. Secondly, the method is iterative, with only one token in the adversarial passage selected and perturbed at each iteration. This process cannot be parallelized within each adversarial passage, as the perturbation of the next token depends on the previously perturbed token. This makes the method time-consuming and resource intensive: in our experiments with a single H100 GPU, HotFlip takes approximately 2 hours to generate a single adversarial passage.
6.2 Performing Corpus Poisoning with Vec2Text
In contrast, Vec2Text does not suffer from the drawbacks of HotFlip. First, Vect2Text does not require access to the embedding model weights. Once Vec2Text is trained, it can directly generate the adversarial passage from the embedding using standard language model generation inference. Second, it allows for efficient generation of a large set of adversarial passages (being far less computationally intensive than HotFlip).
To formally define corpus poisoning with Vec2Text, let $\mathcal{Q}=\{q_{1},q_{2},...,q_{|\mathcal{Q}|}\}$ be a set of training queries. The goal is to use Vec2Text to generate an adversarial passage $a$ whose embedding maximizes the similarity to all the training query embeddings:
$a$ $=\underset{a^{\prime}}{\mathrm{argmax}}\frac{1}{\mathcal{|Q|}}\sum_{q_{i}\in\mathcal{Q}}\phi(q_{i})^{\intercal}\phi(a^{\prime})$ (3)
$=\underset{a^{\prime}}{\mathrm{argmax}}\>\phi(a^{\prime})^{\intercal}\frac{1}{\mathcal{|Q|}}\sum_{q_{i}\in\mathcal{Q}}\phi(q_{i})$ (4)
$=\underset{a^{\prime}}{\mathrm{argmax}}\>\phi(a^{\prime})^{\intercal}\phi_{\mathcal{Q}},$ (5)
where $\phi_{\mathcal{Q}}$ is the average embedding or centroid embedding of all the training query embeddings. The maximum similarity is achieved when $\phi(a^{\prime})=\phi_{\mathcal{Q}}$. Vec2Text can solve this optimization problem through a three-stage process:
  • The first stage is the standard training of Vec2Text (i.e., reconstructing the text for any given embedding).
  • The second stage involves computing the centroid embedding $\phi_{\mathcal{Q}}$ of training queries, which can be done by sending the training queries to the embedding model to obtain all the query embeddings and averaging them.
  • Finally, inputting $\phi_{\mathcal{Q}}$ to the trained Vec2Text yields a generated text, which is the adversarial passage $a$.
A perfect Vec2Text would generate an adversarial passage whose embedding is exactly the same as the query centroid; i.e., $\phi(a)=\phi_{\mathcal{Q}}$.
Once again, we emphasize that the entire process does not require access to model weights. Moreover, generating an adversarial passage with Vec2Text is efficient: generating a single passage took 5 seconds on a single H100 GPU, employing beam search with 50 steps and a width of 4.
Table 7: Results of applying Vec2Text to the task of corpus poisoning, compared to existing HotFlip approach. Vec2Text has the advantage of being able to produce numerous adversarial passages ($k$) without model weights. Both Vec2Text and HotFlip are well below the upper bound, indicating that better methods could pose a serious risk in corpus poisoning.
Attack method #Adversarial Passages / Success@10 Success@20 Success@100 Success@1000
Query Clusters $k$
HotFlip Zhong et al. (2023) 10 0.006 0.015 0.105 0.532
Vec2Text 10 0.006 0.012 0.036 0.101
   Query centroid (upper bound) 10 0.206 0.304 0.575 0.926
HotFlip 100 $\cdots$ Too computationally expensive $\cdots$
Vec2Text 100 0.061 0.091 0.189 0.401
   Query centroid (upper bound) 100 0.493 0.620 0.868 0.994
HotFlip 1000 $\cdots$ Too computationally expensive $\cdots$
Vec2Text 1000 0.268 0.335 0.521 0.791
   Query centroid (upper bound) 1000 0.779 0.873 0.980 1.000
6.3 Corpus Poisoning Experiments
$a$ = "That drawings promises Key kot iner jor Respond machines AK <pad> ance pe izi very nie <pad> <pad> Doar ceapa Weg MEN Am OK Jamie words Um Imp refers ’’ War Hop nism the serving Bol auto Palatul fries rome lighter (1 49 fetch 19 cities can counting 16, As"
(a) HotFlip. Cosine similarity of this passage to the query centroid embedding is $\cos(\phi(a),\phi_{\mathcal{Q}})=0.96$.
$a$ = "when does the 7 episode season of the new two come out </s>"
(b) Vect2Text. Cosine similarity of this passage to the query centroid embedding is $\cos(\phi(a),\phi_{\mathcal{Q}})=0.85$.
Figure 3: Sample adversarial passages most closely matching the NQ training query centroid.
We conducted the corpus poisoning experiments with the GTR-base embedding model (with correct embeddings) as the target DR system, utilizing the NQ datasets as the target corpus. NQ training queries are used to first encode the query centroid embedding and then to generate adversarial passages. In practice, we generate multiple adversarial passages (in line with Zhong et al. (2023)). This was done by first using $k$-means clustering to cluster the training queries based on their embeddings. Once the queries are grouped into clusters, we take the centroid embedding of each cluster and input them to Vec2Text to generate one adversarial passage for each cluster. We set $k$ to $10,100,1000$.
Evaluation was then done using NQ test queries. The evaluation measure is $Success@n$ (in line with Zhong et al. (2023)), which represents the percentage of queries for which at least one adversarial passage is retrieved in the top-$n$ results. A higher success rate indicates a greater vulnerability to corpus poisoning.
Table 7 presents corpus poisoning results. When only 10 adversarial passages are generated (grouping training queries with 10-means clustering), both HotFlip and Vec2Text perform poorly on success@10 and 20; however, HotFlip does work when injecting adversarial passages into the top 100 or 1000 rank positions. If the embedding model weights are available, HotFlip is a more potent corpus poisoning attack method when applied to longer rank lists of 100 or 100.
Vec2Text has the advantage of being able to generate many more adversarial passages; when $k=100$ or $k=1000$ HotFlip is too computationally expensive and thus largely impractical. For these cases Vec2Text success rate improves, with $k=1000$ making Vec2Text a considerable threat in corpus poisoning.
The adversarial passages are generated from query embeddings and are such synthetic and may not appear like real passages. Figure 3 shows the two adversarial passages most closely matching a sampled query centroid $\phi_{\mathcal{Q}}$ for both HotFlip and Vec2Text. The cosine similarity with the query centroid embeddings is also reported in the caption. Recall that HotFlip works by flipping tokens one at a time, whereas Vec2Text is a language model decoder. This difference may explain by Vec2Text appears to generate more natural language.
The embeddings of the adversarial passages in Figure 3 still differ from the query centroid embeddings (cosine similarity being 0.96 for HotFlip and 0.85 for Vec2Text). Recall that a perfect Vec2Text would generate an adversarial passage for which the embedding is exactly the same as the query embedding. To understand how this best case scenario affects success@k we run an additional experiment which involved directly inserting the query centroid embeddings into the vector index and then evaluating the success@k. This serves as an upper bound for the HotFlip and Vec2Text corpus poisoning attacks — an analysis overlooked in the original corpus poisoning paper Zhong et al. (2023). Results are shown in Table 7 as "Query centroid (upper bound)". For both different values of $k$ and different success@$n$, the upper bound is much higher than both HotFlip or Vec2Text. This tells us that neither method is optimal yet for the corpus poisoning attack. It also highlights that if a much more effective corpus poising method — one closer to the upper bound — was developed, it could have serious adverse consequences for dense retrievers.
7 Conclusion
@@
This paper reproduces and extends Vec2Text. We begin by uncovering an issue in the original Vec2Text implementation whereby the last hidden layer rather than the dense pooling layer was used for the final text embeddings for one of the considered embedding models. We correct this issue, rerunning the original experiments and showing that Vec2Text is even more effective in retrieval and reconstructibility.
@@
Finally, we identify that Vec2Text could be applied as a threat to the completely different task for corpus poisoning, whereby adversarial passages are generated and inserted into a corpus such that they are likely be retrieved for any query. Vect2Text poses a great risk here because it can easily generate large numbers of adversarial passages without access to model weights. We show that under certain conditions, corpus poisoning with Vec2Text can pose a serious threat to dense retriever system integrity and user experience.
@@
Acknowledgments and Disclosure of Funding
We thank Morris et al. for giving permission to re-use their figure illustrating the Vec2Text method. This research is partially funded by Beijing Baidu Netcom Technology Co, Ltd, for the project "Federated Online Learning of Neural Rankers", under funding schema 2022 CCF-Baidu Pinecone.

### v2

@@ 4. Understanding what impacts Vec2Text effectiveness
To maintain control over experimental parameters and prevent results from being influenced by factors such as different datasets and training parameters, we independently train the embedding models ourselves. This is done with the same experimental settings but utilizing different training or inference strategies. Specifically, we start by reproducing the DPR embedding model (Karpukhin et al., 2020), following the training parameters outlined in the DPR paper and employing the Tevatron dense retriever training toolkit (Gao et al., 2023) to train the DPR model on the NQ dataset, the same dataset used to train Vec2Text. After reproducing the DPR retrieval results, we then apply the same dense retriever training hyperparameters (e.g., batch size, learning rate, and number of hard negatives) to train other DPR variations with distinct training strategies, as detailed in the next subsections.
Table 4. Zero-shot retrieval and bottlenecked pre-training results. Subscripts denote statistically significantly better retrieval effectiveness with a T-test ($p<0.01$). Results show that even zero-shot retrievers are at risk to Vec2Text reconstruction. Bottleneck pre-training proved useful at improve retrieval effectiveness but greatly increased risk of reconstructibility.
Embedding Models Retrieval effectiveness Reconstructibility
Label Setting top 10 top 20 top 100 top 1000 bleu tf1 exact cos
a. BERT_cls_dot Zero-shot 0.101 0.153 0.288 0.467 47.9 73.3 11.7 0.981
b. BERT_mean_cos Zero-shot 0.283a 0.358a 0.520a 0.715a 86.9 94.1 51.6 0.996
c. SimLM_cls_dot Bottleneck + Zero-shot 0.304ab 0.378a 0.546ab 0.720a 94.4 97.2 64.4 0.999
d. SimLM_cls_dot_finetuned Bottleneck + Fine-tuned 0.750abc 0.805abc 0.874abc 0.919abc 94.7 97.2 63.1 1.000
4.1. Distance Metric and Pooling Method
@@
Regarding reconstructibility, a comparison between DRP_cls_dot and DRP_cls_cos suggests that cosine similarity appears to enhance Vec2Text’s scores. However, when comparing DRP_cls_cos and DRP_mean_cos, the reconstructibility of Vec2Text significantly increases. This surprising result suggests that mean pooling is a key factor for Vec2Text reconstructibility but does not necessarily contribute to improved retrieval effectiveness. We suspect that this might be due to mean pooling explicitly gathering all the token information into the embedding, whereas the CLS token embedding only implicitly learns the text information during DR training, making it less helpful for Vec2Text training. Consequently, practitioners in the field of DR should carefully consider avoiding the use of mean pooling to enhance the security of the embedding model against Vec2Text attacks.
@@
4.2. Zero-shot Regime and Bottleneck Pre-training
@@
BERT_cls_dot and BERT_mean_cos use the same backbone pre-trained embedding model (BERT (Kenton and Toutanova, 2019)) and perform the exact same inference as DRP_cls_dot and DRP_mean_cos but without DPR training; they represent zero-shot retrievers. Their zero-shot retrieval effectiveness is significantly worse than fine-tuned DPR models (as might be expected for a zero-shot dense retriever). Mean pooling with cosine similarity exhibits higher zero-shot retrieval effectiveness than dot product with CLS token embeddings but still well below the trained DPR models from the previous section.
@@
We then tested the SimLM (Wang et al., 2023) embedding model, which leverages a bottlenecked pre-training approach (Gao and Callan, 2021; Lu et al., 2021; Liu and Shao, 2022; Shen et al., 2022; Chuang et al., 2022; Wu et al., 2022; Zhuang et al., 2023). Bottlenecked pre-training is a task designed to enhance the CLS token embedding. This is achieved by taking the embedding model’s outputted CLS token embedding and inputting it into a weaker decoder model. The pre-training objective is to enable the decoder model to perform the MASK language modelling task or, in the case of SimLM, the ELECTRA pre-training task (Clark et al., 2019). Consequently, the model must learn to inject useful information into the CLS token embedding to assist the decoder in completing the task. In fact, this pre-training approach closely resembles Vec2Text training itsself, as both involve a generator that takes an embedding as input and attempts to reconstruct a piece of text. Previous studies on bottlenecked pre-training have shown improvements in DR retrieval effectiveness. However, we suspect that this pre-training approach may make the embedding particularly vulnerable to Vec2Text, raising potential privacy concerns.
@@
Table 5. Impact of embedding dimension and quantization. Both dimension reduction and product quantization prove to be simple and effective methods to mitigate Vec2Text reconstruction, without comprising retrieval effectiveness. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot (768) 768 None 61GB 0.748 0.800 0.863 0.914 79.2 90.9 43.0 0.996
b. DPR_cls_dot (256) 256 None 21GB 0.731 0.784 0.855 0.910 28.0 59.9 5.9 0.966
c. DPR_cls_dot (PQ_768) 768 Product 16GB 0.749 0.801 0.862 0.914 2.2 16.4 0.0 0.772
d. DPR_cls_dot (PQ_256) 256 Product 5GB 0.740 0.796 0.864 0.912 2.5 17.3 0.0 0.782
4.3. Embedding Dimensionality & Quantization
@@
Regarding embedding dimensionality, we trained another DPR model with the exact settings as DPR_cls_dot, but adding a dense pooling layer on top to reduce the embedding dimension from 768 to 256. This variant is labeled as DPR_cls_dot (256). For embedding quantization, we applied product quantization (PQ) (Jegou et al., 2010) to DPR_cls_dot. We experimented with two settings: one with the number of sub-vectors set to 768 and another with 256. The number of bits per sub-vector for both settings is set to 8 in both cases, resulting in embedding dimensions of 768 and 256. We label them DPR_cls_dot (PQ_768) and DPR_cls_dot (PQ_256) respectively.
The results presented in Table 5 indicate that dimensionality reduction and product quantization can significantly reduce the dense vector index sizes, as expected. Product quantization demonstrates robust retrieval effectiveness; when the sub-vector is set to 768, there is no decrease in top-k accuracy, and only a slight decrease in retrieval effectiveness when further reducing the dimensionality to 256. These findings align with previous studies (Ma et al., 2021).
@@
Figure 2. Impact on retrieval effectiveness and reconstructibility of different amounts of noise injection with different retrieval models. Larger $\lambda$ signifies more noise injection.
Table 6. Impact of adding noise and customized embedding transformation. Noise injection was an effective reconstruction mitigation strategy. The proposed vector transformation mitigation strategy is easily applied and guaranteed to maintain retrieval effective while completely degrading reconstructibility. There is no statistically significant difference observed in terms of retrieval effectiveness.
@@
a. DPR_cls_dot None 0.748 0.800 0.863 0.914 85.3 93.8 54.6 0.998
b. DPR_cls_dot_noise Noise injection 0.730 0.790 0.856 0.911 12.7 45.5 0.0 0.879
c. DPR_cls_dot_transform Vector transformation 0.748 0.800 0.863 0.914 1.1 6.0 0.0 0.521
Table 7. Results of adding noise during Text2Text training. The target embedding model is $DPR\_cls\_dot$. Noise injection proved to be an effective mitigation strategy even when Vec2Text is actually trained with noisy embeddings.
@@
Vec2Text Original embeddings 85.3 93.8 54.6 0.998 12.7 45.5 0.0 0.879
Vec2Text_noise Noisy embeddings 27.2 62.6 1.3 0.979 21.0 55.3 0.7 0.908
5. Mitigation Strategies
5.1. Noise Injection
(1) $\phi_{noisy}(x)=\phi(x)+\lambda\cdot\epsilon,\epsilon\sim\mathcal{N}(0,1),$
@@ 5.2. Embedding Transformation
6. Conclusion
@@
This paper reproduces and extends Vec2Text. We begin by uncovering two issues in the original Vec2Text experimental setting. One issue is related to the Vec2Text implementation, whereby the last hidden layer rather than the dense pooling layer was used for the final text embeddings for one of the considered embedding models. The other issue is a serious data leakage problem, where more than 50% of validation data points also appear in the Vec2Text training data. We correct these issues and re-executed the original experiments, providing a more accurate evaluation.
@@
Acknowledgements.
We thank Morris et al. for giving permission to reuse their figure illustrating the Vec2Text method. We also thank Xin Cheng from Peking University for reviewing our code and identifying the data leakage issue in the original NQ dataset used for training Vec2Text models. This research is partially funded by Beijing Baidu Netcom Technology Co, Ltd, for the project "Federated Online Learning of Neural Rankers", under funding schema 2022 CCF-Baidu Pinecone.